Matt makes a cardboard frame to go around the edges of an 8cm * 10cm photograph. The frame has an area of 63 square centimeters and is of equal width all the way around. What is the width?
Answer by edjones(8007)

L*w=A
10*8=80
Let x be the increase in length or width in the frame compared to the picture.
(x+10)(x+8)=63+80
x^2+18x+80=143
x^2+18x-63=0
(x+21)(x-3)=0
x=3
.
Ed